# iCost: A Novel Instance Complexity Based Cost-Sensitive Learning Framework for Imbalanced Classification

### v1

@@ III Methodology
This has been illustrated in Fig. 2. Samples that are categorized as ’pure’ are completely surrounded by instances of the same class. These samples are easy to classify and are usually located far from the decision boundary. Hence, a comparatively much smaller misclassification cost should be enough to correctly identify these samples. Assigning a higher weight might worsen the scenario, resulting in a higher number of misclassifications of the majority class instances. ’Safe’ samples have 1 or 2 neighboring opposite-class instances. These samples should be handled carefully due to the risk of misclassification. Assigning too small a weight may be insufficient, while too large a weight could reverse the situation. On the other hand, border samples are surrounded by majority-class instances. These samples would be misclassified by the K-nearest neighbor classification rule. Therefore, a higher weight is necessary to prioritize these samples over the neighboring samples from the opposite class.
@@
In our implementation, different weights are assigned to different categories of samples based on their difficulty level. We employed a grid-search technique to determine the appropriate cost for different categories. The values varied from one dataset to another. Based on the experiments conducted on 66 datasets, we set a default value for each category for direct implementation. The default setting of the weights also provides quite an improvement in performance. The default penalty value for border samples is fixed as the IR of the dataset. For the safe samples, half of that value is utilized. For the ’pure’ category, a cost factor of 1.2 is selected which is almost the same as the misclassification cost of majority class samples. Employing different search algorithms to find the more suitable weights for individual datasets can optimize the prediction performance.
@@
There are other ways of categorizing minority-class instances. One such way was proposed by Napierala et al. [29] where the samples are classified into four categories: safe, borderline, rare, and outliers. This is also implemented in our proposed framework by assigning four different penalties. Employing search algorithms to find a suitable set of four different values can be a bit more time-consuming. To give the user more freedom, we provide a general categorization formula based on the number of majority-class samples surrounding a minority-class instance. We grade each minority-class sample from g0 to g5 based on 0 to 5 neighboring majority-class samples, respectively. Then, different weights can be assigned to each type of minority-class sample based on its grade.
We implemented our proposed approach using the Python programming language. The theoretical framework behind cost-sensitive classifiers such as CS-SVM has been discussed in detail in previous literature [16] and therefore, not repeated in this manuscript. The implementation code can be found in the following repository: (the GitHub repo could not be released during submission due to blind review policy).
The architecture of our proposed algorithm is presented below:
@@
  • data: Input dataset (Pandas DataFrame)
  • classifier: ’SVM’, ’LR’, or ’DT’ (default = ’SVM’). The algorithm inherits from Sklearn’s SVC, LogisticRegression, and DecisionTreeClassifier implementations, respectively.
  • type: ’org’, ’ins’, ’nap’, or ’gen’ (default = ’ins’). ’org’ refers to the original implementation of the CS classifier. ’ins’ refers to the instance categorization criteria we proposed. ’nap’ refers to the instance categorization criteria proposed by Napierala et al. [29]. ’gen’ provides the general categorization mentioned earlier.
  • k: The number of neighbors to be considered for categorization of the minority-class instances (default = 5).
  • cost-factor: The misclassification cost to be assigned. It can be an integer or a list/dictionary. This input parameter is related to the ’type’ parameter. For type = ’org’, the cost-factor value must be an integer. For other types, the cost-factor value can be both an integer and a list/dictionary. For all cases, the default value is set as the IR of the dataset.
Output: Instance-level weighted classifier fitted on the given input training data.
@@
  • If type = ’org’, the algorithm assigns a weight equal to the cost factor to all the minority-class instances without any other consideration. This is the original CS implementation of the algorithms. If the cost factor value is 1, the algorithm will work as a standard ED classifier.
  • If type = ’ins’ or ’nap’, the algorithm categorizes the minority-class instances into three or four categories, respectively. For ’gen’, minority-class instances are categorized into k+1 categories.
  • In the case of ’ins’, the user can provide an integer or an array/dictionary with three elements as the input values for the cost factor. If it is an integer, then a penalty equal to the integer is assigned to the ’border’ samples. For the ’safe’ samples, half of the integer value is assigned as the penalty factor. For the ’pure’ samples, misclassification cost = 1.2 is fixed. If the input is an array, values are directly assigned to border, safe, and pure samples, in that order. In the case of a dictionary, key-value pairs can be used to directly state the cost values for each pair.
  • Same thing goes for type = ’nap’. If the input value for the cost factor is an integer, then the weights are assigned to the minority class samples in the following way: outlier = cost factor, rare = 0.75 * cost factor, border = 0.5 * cost factor, and safe = 0.25 * cost factor. The user can also directly assign weights using an array or dictionary with four elements.
  • For ’gen’, the user can assign weights using an array of k+1 elements. In the case of integer input or default scenario (weight=IR), the weight is equally divided between the samples from 1 to IR proportionally based on their grade.
  • The appropriate values for the misclassification cost are dependent on the dataset. A value equal to the IR of the dataset is set as default, similar to the implementation of the sklearn library. The above-mentioned approach of assigning costs to different categories of samples provides a decent improvement in performance. However, it can be further optimized using different search algorithms [30].
@@
  • iCost(data, classifier = ’LR’, type = ’gen’, cost-factor = [2, 2, 2, 10, 10, 10])
  • This will apply an instance complexity-based cost-sensitive Logistic Regression (LR) classifier on the given data. Here, g0, g1, and g2 graded minority-class samples (’pure’ and ’safe’ categories) are weighted by a factor of 2. The remaining samples (’border’ category) are weighted by a factor of 10.
IV Experiment
IV-A Datasets
The performance of the proposed algorithm has been evaluated on 66 imbalanced datasets with varying degrees of imbalance to ensure the generalizability of the proposed approach. The datasets are collected from KEEL and the UCI data repository [31]. All the datasets are publicly available and with no missing entries. A summary of the datasets is provided in Table II.
TABLE II: Summary of the datasets
@@ IV-A Datasets
IV-B Experimental Setup
To ensure proper validation and avoid data leakage, the data was first split into training and testing folds. The algorithms are applied only to the training set and the performance is evaluated on the testing set. A repeated stratified K-fold cross-validation strategy (5 folds, 10 repeats) was adopted to ensure better generalization. The average of the results from all different testing folds is considered as the performance measure.
We experimented with 3 different classification algorithms in this study: LR, SVM, and Decision Tree (DT). The default parameters of the Sklearn library were utilized to implement these classifiers. In the case of the proposed iCost algorithm, only the ’cost-factor’ parameter was tuned using the grid-search technique. We used type=’org’ and ’ins’ to obtain results for the traditional CS classifiers and our proposed modified CS approach, respectively. The parameter setting for the grid-search implementation is provided in Table III. MCC score was utilized as the scoring criteria.
TABLE III: Parameter settings for the grid-search implementation of the proposed iCost algorithm
@@ IV-B Experimental Setup
Type org 0.8*IR, IR, 1.2*IR
Type ins ’pure’ : [1, 0.2*IR]
@@
IV-C Performance Metrics
@@
Assessing the performance of different techniques on skewed data can be challenging [1]. Eight different performance metrics were computed for evaluation in this study: MCC, ROC-AUC, G-mean, F1-score, sensitivity, specificity, precision, and accuracy. ML algorithms often excel at predicting instances from the majority class but tend to perform poorly on the minority class. Consequently, traditional performance metrics like accuracy can be misleading because they do not account for the distribution of classes. For instance, if 95% of the data belongs to one class and 5% to another, a model that always predicts the majority class will have 95% accuracy but will be useless for predicting the minority class. Sensitivity and specificity are two class-specific metrics that manifest the performance accuracy of the minority and majority classes, respectively. They only show the performance of a particular class. As a result, it is difficult to apprehend the entire performance spectrum from these measures.
@@
Composite metrics are more suitable for performance measurements on imbalanced data. The g-mean score exhibits a broader picture by providing the geometric mean of sensitivity and specificity. F1-score provides the harmonic mean between sensitivity and precision. The algorithm needs to attain a high sensitivity to have a better G-mean or F1-score. While these two metrics are quite popular and have their advantages, there are certain issues and limitations associated with them. They do not consider the actual number of misclassifications. To elaborate, a sensitivity score of 0.8, observed with only 100 minority-class samples in the dataset, implies that the model misclassified 20 instances from the minority class. On the other hand, a specificity score of 0.8 with 10,000 majority-class samples in the same data indicates that the model misclassified 2,000 instances. This is a huge difference in the number of misclassifications which is not apparent from class-specific metrics. An algorithm may improve the sensitivity score to 0.9 which indicates that only 10 minority class instances have been misclassified. However, improving sensitivity often comes at the cost of lowering the specificity score. If the specificity score is reduced to 0.7 by the algorithm, then the model is now making 3000 misclassifications on the majority class. For both cases, the overall g-mean score remains almost the same. However, the model is now misclassifying around 1000 more samples which is not captured by these metrics.
@@
To be more vigilant regarding performance measurements, we consider other metrics such as MCC and ROC-AUC. MCC provides a robust performance measure as it considers the actual number of misclassifications for both classes. This comprehensive consideration provides a balanced measure of model performance. The MCC score is high only when the classifier performs well across all cases, making it one of the most robust measures of classification performance. However, as stated in several previous literature [32, 33], the performance of algorithms on imbalanced data cannot be sufficiently represented by a single metric. Therefore, we take into consideration four composite metrics to understand how the proposed approach fares compared to other techniques used in imbalanced learning.
IV-D Performance Comparison

### v2

@@ 3.2.1 Neighborhood Search
This has been illustrated in Fig. 4. Samples that are categorized as ’pure’ are completely surrounded by instances of the same class. These samples are easy to classify and are usually located far from the decision boundary. ’Safe’ samples have 1 or 2 neighboring opposite-class instances. These samples should be handled carefully due to the risk of misclassification. Border samples are surrounded by majority-class instances. These samples would be misclassified by the K-nearest neighbor classification rule.
This categorization formula is flexible and can be easily modified. We keep that option in the implementation of the algorithm.
@@
3.2.2 Minimum Spanning Tree (MST)
A spanning tree is a tree-like subgraph of a connected, undirected graph that contains all the vertices of the graph. A minimum spanning tree shares the characteristics of a spanning tree but with the additional condition of having the smallest total edge weights among all possible spanning trees. In this approach, the MST is first built on the dataset by connecting all the instances. Next, all the vertices (data points) that have an edge connecting them from 2 distinct classes are identified. The minority-class instances that are connected to the opposite class by an edge of the MST are penalized heavily compared to other minority-class samples.
This has been illustrated in Fig. 5. In this approach, the minority-class instances are divided into only two categories. While this offers less flexibility, it is easily tunable and provides better performance than the traditional CS approach. Here, unlike the neighborhood approach, the algorithm focuses on the immediate neighbor and penalizes accordingly. The Scipy library was utilized to calculate the MST.
Figure 5: Categorization of Minority-class instances based on MST. The marked points represent linked examples.
3.3 Proposed Algorithm
The proposed approach can be broadly divided into two segments: categorization of the minority-class instances and their weighting.
3.3.1 Minority-class instance categorization
In the first stage. the minority-class instances are categorized based on their complexity. Two criteria have been implemented so far, as described in the previous section.
In the neighborhood search approach, instances are categorized into three 3 distinct groups. For the ’pure’ group, a comparatively much smaller misclassification cost should be enough to correctly identify these samples. Assigning a higher weight might worsen the scenario by unusually deforming the decision boundary. This may result in a higher number of misclassifications of any nearby majority-class instances. For the ’safe’ group, assigning too small a weight may be insufficient, while too large a weight could reverse the situation. On the other hand, for the ’border’ group, a higher weight is necessary to prioritize these samples over the neighboring samples from the opposite class.
In the MST approach, instances are categorized into two distinct groups. The instances that are linked to the opposite class are strongly penalized. These instances are usually located along the border. The remaining instances are marginally penalized. The weight of majority class misclassification is usually kept as 1 for simplicity. However, for the linked majority-class instances, the misclassification cost can be reduced to minimize their effect.
3.3.2 Instance Weighting
@@
In our implementation, different weights are assigned to different categories of samples based on their difficulty level. We employed a grid-search technique to determine the appropriate cost for different categories during experiments. The values varied from one dataset to another. Based on the experiments conducted on 66 datasets, we set a default value for each category for direct implementation. The default setting of the weights also provides quite an improvement in performance. Employing different search algorithms such as evolutionary techniques to find the more suitable weights for individual datasets can further optimize the prediction performance.
The default penalty value for border samples is fixed as the IR of the dataset. For the safe samples, half of that value is utilized. For the ’pure’ category, a cost factor of 0.2*IR is selected which is usually closer to the misclassification cost of majority class samples.
The default penalty value for linked instances is set to the IR of the dataset. As for the other samples, half of IR is used as the misclassification cost.
3.3.3 Additional Features
There are other ways of categorizing minority-class instances. For instance, we can include another category ’outlier’ for those instances that are completely surrounded by opposite classes. The associated penalty for these instances should be limited to avoid overfitting. This can also implemented in our proposed framework by assigning four different penalties. However, employing search algorithms to find a suitable set of four different values can be a bit more time-consuming. While this was not tested in the experiment, we hypothesize that it could prove highly effective in large datasets with many outliers.
@@ 3.3.3 Additional Features
To give the user this freedom, we provide a general categorization formula based on the number of majority-class samples surrounding a minority-class instance. We grade each minority-class sample from g0 to g5 based on 0 to 5 neighboring majority-class samples, respectively. Then, different weights can be assigned to each type of minority-class sample based on its grade. It also opens for some other considerations, such as the presence of outliers. The minority-class samples that are labeled as ’g5’ are completely surrounded by instances of the opposite class. One possible explanation for this could be that these are noisy instances within the majority class. Assigning higher weight to such instances may lead to the misclassification of several neighboring majority-class samples. In the proposed methodology, we have the option to ignore such instances ensuring better generalization.
3.4 Implementation
The proposed approach is implemented using the Python programming language. It is built on top of the scikit-learn library. It inherits from two parent classes: BaseEstimator and ClassifierMixin, to ensure compatibility with scikit-learn’s tools, allowing it to integrate into ML pipelines, perform hyperparameter tuning, and use standard methods like fit, predict, and score.
The architecture is presented below:
@@ 3.4 Implementation
  • 1.
    data: Input dataset (Pandas DataFrame object)
  • 2.
    base_classifier: A classifier instance (e.g., LogisticRegression(), SVC(), RandomForestClassifier(), etc.). The framework inherits from sklearn’s implementations of those algorithms.
  • 3.
    algorithm: ’original’, ’neighborhood’ (default), or ’MST’. Here, ’original’ refers to the traditional implementation of the CS classifiers where all the minority-class instances are weighted equally.
  • 4.
    type: ’ins’ or ’gen’ (default = ’ins’).
    • (a)
      ’ins’ refers to the instance categorization criteria (3 groups) we proposed.
    • (b)
      ’gen’ provides the general categorization (g0 to g6) mentioned earlier.
    These are only applicable if the ’algorithm’ parameter is set to ’neighborhood’.
  • 5.
    n_neighbors: The number of neighbors to be considered for categorization of the minority-class instances under ’neighborhood’ criteria. The default value is set to 5 throughout the experiment.
  • 6.
    cost-factor: The misclassification costs to be assigned. It can be an integer, a list, or a dictionary. This input parameter is related to the ’algorithm’ and ’type’ parameters.
    • (a)
      For algorithm = ’original’, the cost-factor value must be an integer. The default is set to the IR of the dataset.
    • (b)
      For algorithm = ’neighborhood’ and type = ’ins’, there are three different weight parameters: cfb, cfs, and cfp. These refer to the costs associated with the border, safe, and pure samples, respectively. Evidently, $cfb>cfs>cfp$. The default values are set as IR, 0.5*IR, and 0.25*IR, respectively.
    • (c)
      For algorithm = ’MST’, there are two different weight parameters: cfl, and cfn. These refer to the costs associated with the linked and normal samples, respectively. The default values are set as IR and 0.5*IR, respectively.
Output: Instance-level weighted classifier object.
Constraints: The base classifier must support the ”sample_weight” parameter. This enables the classifiers to be made cost-sensitive. Certain classification algorithms (Naive Bayes, for instance) are not inherently designed for cost-sensitive classification. So, they cannot be used as a base classifier.
@@
  • 1.
    If algorithm = ’original’, the algorithm assigns a weight equal to the cost factor to all the minority-class instances without any other consideration. This is the original CS implementation of the algorithms. If the cost factor value is 1, the algorithm will work as a standard ED classifier.
  • 2.
    If type = ’ins’, the algorithm categorizes the minority-class instances into three categories. For ’gen’, minority-class instances are categorized into k+1 categories.
  • 3.
    In the case of ’ins’, the user can provide an integer, an array, or a dictionary with three elements as the input values for the cost factor. If it is an integer, then the following penalty values are assigned.
    • (a)
      ’border’ samples: a penalty equal to the integer is assigned.
    • (b)
      ’safe’ samples: half of the integer value is assigned.
    • (c)
      ’pure’ samples: 0.25 times the integer value is assigned.
    If the input is an array, values are directly assigned to border, safe, and pure samples, in that order. In the case of a dictionary, key-value pairs can be used to directly state the cost values for each pair.
    • (a)
      list: cost-factor = [10, 4, 1.5]
    • (b)
      dictionary: cost-factor = {’cfb’: 10, ’cfs’: 4, ’cfp’: 1.5}
  • 4.
    The same thing goes for algorithm = ’MST’. Here, the minority-class instances are categorized into two distinct groups: linked and normal.
    If the input value for the cost factor is an integer, then the weights are assigned to the minority class samples in the following way:
    • (a)
      linked = cost factor
    • (b)
      normal = 0.5 * cost factor
    The user can also directly assign weights using an array or dictionary with two elements.
  • 5.
    For ’gen’, the user can assign weights using an array of k+1 elements. In the case of integer input or default scenario (weight=IR), the weight is equally divided between the samples from 1 to IR proportionally based on their grade. A dictionary can also be utilized with keys being grades g0 to g6.
Other Considerations: There are a few other issues that should be considered while applying the algorithm. They are discussed below.
  • 1.
    The appropriate values for the misclassification cost are dependent on the dataset. In the implementation of the sklearn library, a penalty equal to the IR of the dataset is employed. However, that value does not necessarily produce the optimal result. We performed experiments with different cost values on several imbalanced datasets and observed that the optimal value usually lies close to the IR of the dataset.
    In the proposed algorithm, the minority-class instances are divided into three groups which require three different cost values. Experiments were conducted with different cost settings on several imbalanced datasets and based on the observation, the default values are set for border, safe, and pure category samples. These default cost settings provide a decent improvement in performance over the traditional CS classifiers.
    However, the performance can be further optimized using different search algorithms such as grid search, random search, bayesian optimization, evolutionary algorithms, etc. [42]. The cost factor is the most sensitive parameter in our proposed framework and we suggest tuning it for optimal results.
  • 2.
@@
  • 1.
    iCost(base_classifier = LogisticRegression(), algorithm = ’neighborhood’, type = ’gen’, cost-factor = [2, 5, 5, 10, 10, 1])
  • 2.
    This will apply an instance complexity-based cost-sensitive LR classifier on the given data. Here,
    • (a)
      g0 graded minority-class samples (’pure’ category) are weighted by a factor of 2.
    • (b)
      The g1 and g2 graded minority-class samples (’safe’ category) are weighted by a factor of 5.
    • (c)
      The g3 and g4 graded samples are weighted by a factor of 10.
    • (d)
      The remaining samples (g6 category) which are entirely surrounded by 5 opposite-class instances are more likely to be outliers and penalizing such instances highly can cause overfitting on minority-class data. Therefore, those are weighted by a factor of 1.
  • 3.
    While such intuitive settings of the cost factor parameter are likely to improve the prediction performance, it is better to tune the parameter to find the most appropriate cost settings and obtain optimal results.
3.5 Extension - Multiclass Classification
The proposed algorithm can be applied for multiclass classification as well. Compared to binary classification scenarios, multiclass classification is much more complicated and requires more sophisticated strategies [43].
Classification through binarization i.e. decomposition strategies such as One-vs-One (OvO) or One-vs-Rest (OvR) are the most commonly used approach used in multiclass classification. In these approaches, the multiclass problem is disintegrated into multiple binary classification problems. These methods have their own drawbacks. For instance, in OvR, even when the classes are balanced, the binarization technique forces the data to become skewed, making the classification more complex [44]. However, these techniques are quite effective in handling data in multiple classes.
The sklearn library supports such decomposition techniques through classes such as OneVsRestClassifier and OneVsOneClassifier. These classifiers take an estimator object as the input. We also ensured the compatibility of our proposed framework with these approaches. To apply the proposed framework in multiclass scenarios, the iCost object just needs to be fed as the estimator to those classifiers. For the experiment, we utilized the OneVsRestClassifier with the 4 base classifiers as estimators to obtain the results.
4 Experiment
4.1 Datasets
The performance of the proposed algorithm has been evaluated on datasets with varying degrees of imbalance to ensure the generalizability of the proposed approach. The datasets are collected from KEEL and the UCI data repository [45]. All the datasets are publicly available. A handful of them have a few missing entries. The missing values are imputed using the MICE algorithm [46]. A total of 65 binary classification datasets and 10 multiclass classification datasets were utilized in this study. A summary of the datasets is provided in Table 2 and Table 3.
Table 2: Summary of the datasets for binary classification tasks
@@ 4.1 Datasets
Table 3: Summary of the datasets for multiclass classification tasks
Dataset # Features # Classes No of Samples/Class
thyroid 21 3 $6$, 368, 166
new-thyroid 5 3 150, 35, 30
contraceptive 9 3 629, 511, 333
shuttle 9 3 1706, 338, 131
Dry_Bean_Dataset 16 7 3546, 2636, 2027, 1928, 1630, 1322, 522
balance 4 3 288, 288, 49
pageblocks 10 5 4913, 329, 115, 88, 28
HCV 12 4 540, 30, 24, 21
yeast 8 10 463, 429, 244, 163, 51, 44, 35, 30, 20, 5
wine 13 3 71, 59, 48
4.2 Experimental Setup
To ensure proper validation and avoid data leakage, the data was first split into training and testing folds. The algorithms are applied only to the training set and the performance is evaluated on the testing set. A repeated stratified K-fold cross-validation strategy (5 folds, 10 repeats) was adopted to ensure better generalization. The average of the results from all different testing folds (50) is considered as the performance measure.
Before training, the data was normalized using z-score (StandardScaler from sklearn library). We experimented with 4 different classification algorithms in this study: LR, SVM, DT, and Random Forest (RF). The default parameters of the scikit-learn library were utilized to implement these classifiers. None of the parameters (except ’class_weight) of the base classifiers were tuned to ensure a valid comparison.
In the case of the proposed iCost algorithm, only the ’cost-factor’ parameter was tuned using the grid-search technique. We used type = ’ins’ to obtain results for the proposed modified CS approach. The parameter setting for the grid-search implementation is provided in Table 4. The MCC score was utilized as the scoring criteria. The performance from other different types of categorization available in the framework may vary slightly. However, those are not included in this manuscript.
Table 4: Parameter settings for the grid-search implementation
@@ 4.2 Experimental Setup
Algorithm original 0.8*IR, 0.9*IR, IR, 1.1*IR, 1.2*IR
Algorithm Neighborhood ’pure’ : [1, 0.2*IR]
@@
Algorithm MST ’linked’ : [0.75*IR, 0.9*IR, IR, 1.1*IR, 1.25*IR]
’normal’: [0.3*IR, 0.5*IR, 0.7*IR]
4.3 Performance Metrics
@@
Assessing the performance of different techniques on skewed data can be challenging [1]. Eight different performance metrics were computed for evaluation in this study: MCC, ROC-AUC, G-mean, F1-score, sensitivity, specificity, precision, and accuracy. ML algorithms often excel at predicting instances from the majority class but tend to perform poorly on the minority class. Consequently, traditional performance metrics like accuracy can be misleading because they do not account for the distribution of classes.
To give an example, if 95% of the data belongs to one class and 5% to another, a model that always predicts the majority class will have 95% accuracy but will be useless for predicting the minority class. Sensitivity and specificity are two class-specific metrics that manifest the performance accuracy of the minority and majority classes, respectively. They only show the performance of a particular class. As a result, it is difficult to apprehend the entire performance spectrum from these measures.
@@
Composite metrics are more suitable for performance measurements on imbalanced data [47, 48]. The g-mean score exhibits a broader picture by providing the geometric mean of sensitivity and specificity. F1-score provides the harmonic mean between sensitivity and precision. The algorithm needs to attain a high sensitivity to have a better G-mean or F1-score. While these two metrics are quite popular and have their advantages, there are certain issues and limitations associated with them. They do not consider the actual number of misclassifications.
To elaborate, a sensitivity score of 0.8, observed with only 100 minority-class samples in the dataset, implies that the model misclassified 20 instances from the minority class. On the other hand, a specificity score of 0.8 with 10,000 majority-class samples in the same data indicates that the model misclassified 2,000 instances. This is a huge difference in the number of misclassifications which is not apparent from class-specific metrics. An algorithm may improve the sensitivity score to 0.9 which indicates that only 10 minority class instances have been misclassified. However, improving sensitivity often comes at the cost of lowering the specificity score. If the specificity score is reduced to 0.7 by the algorithm, then the model is now making 3000 misclassifications on the majority class. For both cases, the overall g-mean score remains almost the same. However, the model is now misclassifying around 1000 more samples which is not captured by these metrics.
@@
To be more vigilant regarding performance measurements, we consider other metrics such as MCC and ROC-AUC [49]. MCC provides a robust performance measure as it considers the actual number of misclassifications for both classes [50]. This comprehensive consideration provides a balanced measure of model performance. The MCC score is high only when the classifier performs well across all cases, making it one of the most robust measures of classification performance [51]. However, as stated in several previous literature [47, 52], the performance of algorithms on imbalanced data cannot be sufficiently represented by a single metric. Therefore, we take into consideration four composite metrics to understand how the proposed approach fares compared to other techniques used in imbalanced learning.
For multiclass settings, three different averaging criteria are available: macro, micro, and weighted [53]. The macro average is usually considered the best for imbalanced datasets as it will give a lower score if the model performs poorly on less frequent classes, even if the majority class is well-predicted. Therefore, the macro-average results are calculated and presented in this manuscript.
4.4 Performance Comparison
One of the main objectives of the developed methodology is to achieve better generalization and performance over the traditional CS implementations of different ML algorithms. As such, we compared the performance of the proposed approach with the traditional CS learning techniques to evaluate the differences.